OpenStudy (okdutchman7):

Please help. WILL MEDAL! @Erak @jim_thompson5910 @Mehek14 @triciaal @hartnn @sleepyjess @johnweldon1993 @satellite73 @dumbcrow

OpenStudy (okdutchman7):

\[2 + \sqrt{3} \csc \theta = 0\]

OpenStudy (erak):

isolate for csc theta

OpenStudy (okdutchman7):

\[0 \le \theta < 360\]

OpenStudy (okdutchman7):

\[csc \theta = \frac{ -2 }{ \sqrt{3} }\]

OpenStudy (erak):

yeah, so do you know what the reciprocal of csc is?

OpenStudy (okdutchman7):

sine function

OpenStudy (okdutchman7):

So what do I do next?

OpenStudy (erak):

So in order to get sin, it's 1/csc right? Get the reciprocal of the other side as well

OpenStudy (okdutchman7):

\[\frac{ 1 }{ \sin } \theta = - \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (erak):

it'd be just sin theta, though. Do you know how to use the unit circle/CAST rule to solve it afterwards?

OpenStudy (okdutchman7):

No

jimthompson5910 (jim_thompson5910):

@okdutchman7 you should have this as part of your steps \[\Large \csc(\theta) = -\frac{2}{\sqrt{3}}\] \[\Large \frac{1}{\sin(\theta)} = -\frac{2}{\sqrt{3}}\] \[\Large \frac{\sin(\theta)}{1} = -\frac{\sqrt{3}}{2}\] \[\Large \sin(\theta) = -\frac{\sqrt{3}}{2}\]

OpenStudy (okdutchman7):

okay so i flipped it too early.

OpenStudy (okdutchman7):

Thank You! the answer is 240 and 300.

jimthompson5910 (jim_thompson5910):

`the answer is 240 and 300. ` yes those are the two solutions for theta where 0 <= theta < 360